OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 4) with a slope of –1 in standard form. –x + 4y = –1 x – 4y = –1 x + y = 3 x – y = –3

OpenStudy (jdoe0001):

can you follow this example? http://www.teachhandson.com/calculationb.gif

OpenStudy (jdoe0001):

that is,using the point-slope intercept form

OpenStudy (compassionate):

Hello, Pam. Very beautiful name. Do you know the formula for putting an equation in Ax + Bx = Y, given the point-slope formula?

OpenStudy (anonymous):

it's a timed exam so I just need an answer please.

OpenStudy (anonymous):

Thank you, @Compassionate

OpenStudy (compassionate):

Write the equation of the line that passes through (–1, 4) with a slope of –1 in standard form. You want to use the formula: y - y1 = m(x - x1) y1 = 4 x1 = -1 m = -1 Plug 'em in: y - 4 = -1(x - (1-)) y - 4 = -1(x + 1) y - 4 = -x - 1 y = 4/x + 4 Now put that in Ax + Bx + = Y I cannot do your work any further from this point. ~~